# MOS-Attack: A Scalable Multi-objective Adversarial Attack Framework

### v2

@@ Abstract
Crafting adversarial examples is crucial for evaluating and enhancing the robustness of Deep Neural Networks (DNNs), presenting a challenge equivalent to maximizing a non-differentiable 0-1 loss function. However, existing single objective methods, namely adversarial attacks focus on a surrogate loss function, do not fully harness the benefits of engaging multiple loss functions, as a result of insufficient understanding of their synergistic and conflicting nature. To overcome these limitations, we propose the Multi-Objective Set-based Attack (MOS Attack), a novel adversarial attack framework leveraging multiple loss functions and automatically uncovering their interrelations. The MOS Attack adopts a set-based multi-objective optimization strategy, enabling the incorporation of numerous loss functions without additional parameters. It also automatically mines synergistic patterns among various losses, facilitating the generation of potent adversarial attacks with fewer objectives. Extensive experiments have shown that our MOS Attack outperforms single-objective attacks. Furthermore, by harnessing the identified synergistic patterns, MOS Attack continues to show superior results with a reduced number of loss functions.
@@ 1 Introduction
Deep neural network (DNN) models have significantly advanced the field of computer vision [26, 38, 15, 28], yet they are vulnerable to adversarial examples [43, 20]. Such examples are inputs that have been subtly modified to cause misclassification, potentially leading to catastrophic consequences in real-world scenarios [16, 7, 19]. Consequently, the development of sophisticated adversarial attack algorithms is crucial for evaluating and enhancing the robustness of these models [11, 34, 52]. However, devising these algorithms presents inherent challenges due to the non-differentiable nature of the original optimization problem, necessitating the use of surrogate loss functions [21] to facilitate gradient-based adversarial attacks [20, 34, 11].
The metric for measuring misclassification is the non-differentiable 0-1 loss function, which surrogate loss functions endeavor to approximate [30]. Adversarial attacks are designed to generate a perturbation $\bm{\delta}$ that causes the misclassification of an input ${\bm{x}}$ with its corresponding label ${\bm{y}}$. This can be formulated as [34, 21]:
@@
Considering the computational intractability of the problem in Equation (1[3], contemporary research commonly employs a differentiable surrogate loss function in place of the 0-1 loss function. This approach enables the utilization of gradient-based optimization techniques to address the resultant surrogate optimization problem. It has propelled considerable progress in gradient-based algorithms, including the Fast Gradient Sign Method (FGSM) [20], Projected Gradient Descent (PGD) [34], and Carlini & Wagner (C&W) attack [8]. Notably, the versatility of the PGD attack has increased with the adoption of diverse surrogate loss functions (APGD-CE, APGD-DLR) [41] and the integration of sophisticated optimization techniques (ACG) [51]. These developments have given rise to more sophisticated adversarial methods.
While single-objective attacks have attracted considerable attention, there is an emerging trend towards integrating multiple loss functions to bolster the attack’s efficacy. Some early endeavors include using multiple targeted loss functions to guide untargeted attack [21] and the strategic alternation of loss functions in the attack process[2]. Furthermore, the adoption of diverse surrogate loss functions such as GAMA [41], BCE [46], and DLR [11] has been instrumental in advancing adversarial attacks.
@@
We have implemented our approach using four widely recognized surrogate loss functions as outlined in previous research [41, 46], as well as four additional functions identified through extensive loss function searches [30, 50]. The resulting MOS-8111We use MOS-8 Attack to denote MOS Attack implemented with 8 loss functions, MOS-3 to denote the attack implemented with three selected loss functions. Attack has proven highly effective through extensive experimentation on the CIFAR-10 [28] and ImageNet [15] datasets, outperforming state-of-the-art methods that leverage advanced gradient-based optimization or eight distinct single-objective attacks for each surrogate loss. Moreover, by examining the synergistic patterns uncovered by MOS-8 Attack, we have developed a tri-objective attack, MOS-3, which has also shown superior performance.
@@ 2 Background
Adversarial attacks encompass methods that create adversarial examples, which are used to assess and enhance model robustness [34, 11]. A white-box threat model is often considered for evaluating adversarial robustness, where the adversary has full access to the model’s architecture, parameters, and gradients. While white-box existing strategies mainly focus on one surrogate loss function [1, 25, 18, 51], a recent trend is the integration of multiple loss functions into the attack paradigm [44, 14, 42, 5, 33].
Single-Objective Attacks. White-box attack methodologies typically employ a singular surrogate loss function, focusing on optimization to craft adversarial examples. Established strategies include the FGSM [20], C&W attack [8], and PGD attack [34]. Croce et al. proposed a novel parameter-free approach, Auto-PGD (APGD) attack, utilizing both Cross Entropy (CE) and the Difference of Logits Ratio (DLR) loss functions. These were subsequently incorporated into the AutoAttack framework as APGD-CE and APGD-DLR [11]. Expanding upon this, Yamamura et al. enhanced APGD with conjugate gradient techniques, resulting in the creation of the powerful Auto Conjugate Gradient (ACG) attack [51].
@@
Multi-Objective Attacks. Recent advancements in adversarial research have involved the integration of multiple surrogate loss functions into the attack framework. Gowal et al. introduced multiple targeted losses to enhance untargeted PGD attacks [21]. Further, work by Nikolaos et al. established that the strategic variation of surrogate loss functions considerably improve adversarial attack performance [2]. However, these studies typically lack a systematic approach and a solid theoretical underpinning for managing multiple losses.
Concurrently, researchers have expanded the adversarial attack framework by introducing other types of objectives. Williams et al. investigated the inclusion of additional norm constraints [48], while Guo et al. and Liu et al. have investigated the trade-off between perturbation intensity and confidence measures [24, 33]. These efforts have contributed to the development of more diversified attack methodologies.
@@ 3.1 Multi-Objective Adversarial Attack
The notation remains consistent with the single-objective scenario as depicted in Equation (1). Furthermore, this paper adopts the extensively utilized $\ell_{\infty}$-ball as the constraint set for perturbations, denoted by =$\mathcal{B}=\{\bm{\delta}:\|\bm{\delta}\|_{\infty}\leq\epsilon\}$.
Examples with higher values across multiple surrogate loss functions are more susceptible to misclassification by the model. In existing literature, this statement is supported by frequent misclassifications of adversarial examples with high values on singular loss functions [34, 11, 51].
@@ 3.2 Decomposition-Based Optimization
In this research, we employ the Tchebycheff decomposition to transform the multi-objective problem into a suite of single-objective subproblems. Contrary to the linear scalarization method, the Tchebycheff approach is capable of targeting any location on the Pareto front. This is well-recognized in the discipline of multi-objective optimization [13, 53, 35]. By adopting this method, given $K$ weight vectors ${\bm{w}}_{1},\ldots,{\bm{w}}_{K}$, the $k$-th decomposed subproblem is defined as follows:
@@
The Tchebycheff method is instrumental in identifying both convex and non-convex parts of the Pareto front with vertical contour lines [45], as demonstrated in Figure 1(a). Nonetheless, it presents three challenges in optimization:
  • Complexity: Accurate approximation of the Pareto front necessitates multiple points, exceeding the number of objectives ($>m$).
@@ 3.3 Smooth Set-based Optimization
Smooth Set-based Optimization To smooth the above optimization problem, we need to take advantage of smooth max and smooth min operators [6, 32, 31].
@@
where $\mu$ is a smoothing parameter. A proof of the above approximation can be found in [6].
@@
The above formulation eases the optimization process and avoids oscillations in the optimization process, which is analyzed in the multi-objective literature [32]. We illustrate the smoothed set-based optimization problem along with a possible optimization trajectory in Figure 1(c).
@@ 4.1 MOS Attack: Implementation by APGD
By formulating a smooth set-based optimization problem, we can now apply gradient-based optimization algorithms for its efficient resolution. Within the domain of adversarial attacks, our framework incorporates the well-known APGD algorithm [11]. In this section, we provide a detailed explanation of our attack, as outlined in Algorithm 1.
@@
0 Cross Entropy Loss [43, 34, 20, 29, 8] $-{\bm{h}}_{y}({\bm{x}})+\log(\sum_{i=1}^{K}e^{{\bm{h}}_{i}({\bm{x}})})$
1 Marginal Loss [10, 8, 21, 11, 41] $-{\bm{h}}_{y}({\bm{x}})+\max_{j\neq y}{\bm{h}}_{j}({\bm{x}})$
2 Difference of Logits Ratio [10] $(-{\bm{h}}_{y}({\bm{x}})+\max_{j\neq y}{\bm{h}}_{j}({\bm{x}}))/({\bm{h}}_{\pi_{1}}({\bm{x}})-{\bm{h}}_{\pi_{3}}({\bm{x}}))$
3 Boosted Cross-Entropy Loss [46] $-\log{\bm{p}}_{y}({\bm{x}})-\log(1-\max_{j\neq y}{\bm{p}}_{j}({\bm{x}}))$
4 Searched Loss 1 [50] $\sum_{i}\exp(10{\bm{p}}/\max_{j}{\bm{p}}_{j})$
5 Searched Loss 2 [50] $\exp(-\max(\text{softmax}({\bm{h}}+2\text{softmax}(5{\bm{h}}))))$
6 Searched Loss 3 [50] $\text{softmax}(-\text{softmax}(2\exp({\bm{h}}){\bm{h}}))(\text{softmax}(2{\bm{h}})+2{\bm{y}}_{\text{one-hot}})$
7 Searched Loss 4 [50] $(\text{softmax}(\text{softmax}(2{\bm{h}})+{\bm{h}}-{\bm{y}}_{\text{one-hot}})-{\bm{y}}_{\text{one-hot}})^{2}$
@@ 5 Searched Loss 2 [50] $\exp(-\max(\text{softmax}({\bm{h}}+2\text{softmax}(5{\bm{h}}))))$
3:  $\bm{X}^{(0)}\leftarrow{\bm{x}}+\mathbf{\Delta}^{(0)}$
4:  $\bm{X}^{(1)}\leftarrow P_{\mathcal{B}}(\bm{X}^{(0)}+\eta\nabla g(\mathbf{\Delta}^{(0)}))$
5:  $\mathbf{\Delta}^{(1)}\leftarrow\bm{X}^{(1)}-{\bm{x}}$
6:  $g_{\text{max}}\leftarrow\max\left\{g(\mathbf{\Delta}^{(0)}),g(\mathbf{\Delta}^{(1)})\right\}$
7:  $\bm{X}_{\text{max}}\leftarrow\bm{X}^{(0)}$ if $g_{\text{max}}\equiv g(\mathbf{\Delta}^{(0)})$ else $\bm{X}_{\text{max}}\leftarrow\bm{X}^{(1)}$
@@ 4.2 Automated Synergistic Pattern Mining
Few solutions automatically maximize different loss functions in groups in smooth set-based optimization [31]. To mine these loss synergistic patterns, we propose an automated mining method. This method includes two steps: 1) determining the dominant examples that contribute to the loss maximization, and 2) determining the synergistic pattern of these dominant examples.
@@
CIFAR-10 ($\epsilon=8/255$) APGD (1) APGD (5) ACG (5) All (1)*8 MOS-8 (1) MOS-8 (5) Diff.(5) MOS||||CE
@@
0 Rade et al. (2022) [36] (ddpm) PreActResNet-18 39.17 39.28 42.65 42.78 (6) 42.59 42.77 +3.49
1 Rade et al. (2022) [36] (extra) PreActResNet-18 38.55 38.72 42.12 42.21 (6) 42.03 42.23 +3.51
2 Sehwag et al. (2022) [40] ResNet-18 41.57 41.76 44.79 44.16 (6) 43.79 44.18 +2.42
3 Chen et al. (2020) [9] ResNet-50 45.80 45.95 48.28 48.04 (4) 48.09 48.14 +3.49
4 Gowal et al. (2020) [22] WideResNet-28-10 34.31 34.46 36.90 36.96 (6) 36.77 36.95 +2.19
5 Wang et al. (2023) [47] WideResNet-28-10 29.72 29.91 N/A 32.44 (6) 32.25 32.49 +2.58
6 Rebuffi et al. (2021) [37] WideResNet-28-10 35.97 36.15 38.80 39.05 (6) 38.91 39.14 +2.99
7 Sehwag et al. (2022) [40] WideResNet-34-10 36.85 36.96 40.18 39.36 (5) 38.97 39.38 +2.43
8 Rade et al. (2022) [36] WideResNet-34-10 34.29 34.45 36.83 36.97 (6) 36.69 36.94 +2.49
9 Gowal et al. (2021) [23] WideResNet-70-16 31.43 31.62 33.04 33.50 (5) 33.33 33.51 +1.89
10 Gowal et al. (2020) [22] WideResNet-70-16 31.89 32.07 33.70 33.94 (5) 33.72 33.92 +1.85
11 Rebuffi et al. (2021) [37] WideResNet-70-16 30.45 30.72 32.75 33.06 (6) 32.79 33.10 +2.38
Average Rank 5.92 4.92 2.79 2.00 3.50 1.58
@@
12 Salman et al. (2020) [39] ResNet-18 70.60 70.74 73.72 74.38 (5) 74.24 74.52 +3.87
13 Salman et al. (2020) [39] ResNet-50 61.38 61.58 63.70 64.92 (7) 64.5 64.94 +3.36
14 Wong et al. (2020) [49] ResNet-50 70.28 70.46 71.94 73.20 (5) 72.96 73.10 +2.64
15 Engstrom et al. (2019) [17] ResNet-50 67.62 67.82 68.60 70.12 (5) 69.86 69.92 +2.10
16 Salman et al. (2020) [39] WideResNet-50-2 59.02 59.12 59.92 61.26 (5) 60.76 61.14 +2.02
@@
: reported results from the paper [51], with attack step $N_{\text{iter}}=100$.

### v3

@@ Abstract
Crafting adversarial examples is crucial for evaluating and enhancing the robustness of Deep Neural Networks (DNNs), presenting a challenge equivalent to maximizing a non-differentiable 0-1 loss function. However, existing single objective methods, namely adversarial attacks focus on a surrogate loss function, do not fully harness the benefits of engaging multiple loss functions, as a result of insufficient understanding of their synergistic and conflicting nature. To overcome these limitations, we propose the Multi-Objective Set-based Attack (MOS Attack), a novel adversarial attack framework leveraging multiple loss functions and automatically uncovering their interrelations. The MOS Attack adopts a set-based multi-objective optimization strategy, enabling the incorporation of numerous loss functions without additional parameters. It also automatically mines synergistic patterns among various losses, facilitating the generation of potent adversarial attacks with fewer objectives. Extensive experiments have shown that our MOS Attack outperforms single-objective attacks. Furthermore, by harnessing the identified synergistic patterns, MOS Attack continues to show superior results with a reduced number of loss functions. Our code is available at https://github.com/pgg3/MOS-Attack.
@@ 1 Introduction
Deep neural network (DNN) models have significantly advanced the field of computer vision [he:2016:deep, sabour:2017:dynamic, deng:2009:imagenet, CIFAR10], yet they are vulnerable to adversarial examples [szegedy:2014:intriguing, goodfellow:2015:explaining]. Such examples are inputs that have been subtly modified to cause misclassification, potentially leading to catastrophic consequences in real-world scenarios [deng:2019:efficient, cao:2019:adversarial, laleh:2022:adversarial]. Consequently, the development of sophisticated adversarial attack algorithms is crucial for evaluating and enhancing the robustness of these models [croce:2020:reliable, madry:2018:towards, zhang:2019:theoretically]. However, devising these algorithms presents inherent challenges due to the non-differentiable nature of the original optimization problem, necessitating the use of surrogate loss functions [gowal:2019:alternative] to facilitate gradient-based adversarial attacks [goodfellow:2015:explaining, madry:2018:towards, croce:2020:reliable].
The metric for measuring misclassification is the non-differentiable 0-1 loss function, which surrogate loss functions endeavor to approximate [li:2022:autoloss]. Adversarial attacks are designed to generate a perturbation $\bm{\delta}$ that causes the misclassification of an input $\bm{x}$} with its corresponding label $\bm{y}$}. This can be formulated as [madry:2018:towards, gowal:2019:alternative]:
@@
Considering the computational intractability of the problem in Equation (1[arora:1997:hardness], contemporary research commonly employs a differentiable surrogate loss function in place of the 0-1 loss function. This approach enables the utilization of gradient-based optimization techniques to address the resultant surrogate optimization problem. It has propelled considerable progress in gradient-based algorithms, including the Fast Gradient Sign Method (FGSM) [goodfellow:2015:explaining], Projected Gradient Descent (PGD) [madry:2018:towards], and Carlini & Wagner (C&W) attack [carlini:2017:towards]. Notably, the versatility of the PGD attack has increased with the adoption of diverse surrogate loss functions (APGD-CE, APGD-DLR) [sriramanan:2020:guided] and the integration of sophisticated optimization techniques (ACG) [yamamura:2022:diversified]. These developments have given rise to more sophisticated adversarial methods.
While single-objective attacks have attracted considerable attention, there is an emerging trend towards integrating multiple loss functions to bolster the attack’s efficacy. Some early endeavors include using multiple targeted loss functions to guide untargeted attack [gowal:2019:alternative] and the strategic alternation of loss functions in the attack process[nikolaos:2022:alternating]. Furthermore, the adoption of diverse surrogate loss functions such as GAMA [sriramanan:2020:guided], BCE [wang:2020:improving], and DLR [croce:2020:reliable] has been instrumental in advancing adversarial attacks.
@@
We have implemented our approach using four widely recognized surrogate loss functions as outlined in previous research [sriramanan:2020:guided, wang:2020:improving], as well as four additional functions identified through extensive loss function searches [li:2022:autoloss, xia:2024:tightening]. The resulting MOS-8111We use MOS-8 Attack to denote MOS Attack implemented with 8 loss functions, MOS-3 to denote the attack implemented with three selected loss functions. Attack has proven highly effective through extensive experimentation on the CIFAR-10 [CIFAR10] and ImageNet [deng:2009:imagenet] datasets, outperforming state-of-the-art methods that leverage advanced gradient-based optimization or eight distinct single-objective attacks for each surrogate loss. Moreover, by examining the synergistic patterns uncovered by MOS-8 Attack, we have developed a tri-objective attack, MOS-3, which has also shown superior performance.
@@ 2 Background
Adversarial attacks encompass methods that create adversarial examples, which are used to assess and enhance model robustness [madry:2018:towards, croce:2020:reliable]. A white-box threat model is often considered for evaluating adversarial robustness, where the adversary has full access to the model’s architecture, parameters, and gradients. While white-box existing strategies mainly focus on one surrogate loss function [antoniou:2020:square, guo:2024:lautoda, fu:2022:autoda, yamamura:2022:diversified], a recent trend is the integration of multiple loss functions into the attack paradigm [wang:2024:amulti, deist:2023:multi, Suzuki:2019:adversarial, baia:2021:effective, liu:2024:effective].
Single-Objective Attacks. White-box attack methodologies typically employ a singular surrogate loss function, focusing on optimization to craft adversarial examples. Established strategies include the FGSM [goodfellow:2015:explaining], C&W attack [carlini:2017:towards], and PGD attack [madry:2018:towards]. Croce et al. proposed a novel parameter-free approach, Auto-PGD (APGD) attack, utilizing both Cross Entropy (CE) and the Difference of Logits Ratio (DLR) loss functions. These were subsequently incorporated into the AutoAttack framework as APGD-CE and APGD-DLR [croce:2020:reliable]. Expanding upon this, Yamamura et al. enhanced APGD with conjugate gradient techniques, resulting in the creation of the powerful Auto Conjugate Gradient (ACG) attack [yamamura:2022:diversified].
@@
Multi-Objective Attacks. Recent advancements in adversarial research have involved the integration of multiple surrogate loss functions into the attack framework. Gowal et al. introduced multiple targeted losses to enhance untargeted PGD attacks [gowal:2019:alternative]. Further, work by Nikolaos et al. established that the strategic variation of surrogate loss functions considerably improve adversarial attack performance [nikolaos:2022:alternating]. However, these studies typically lack a systematic approach and a solid theoretical underpinning for managing multiple losses.
Concurrently, researchers have expanded the adversarial attack framework by introducing other types of objectives. Williams et al. investigated the inclusion of additional norm constraints [williams:2023:black-box], while Guo et al. and Liu et al. have investigated the trade-off between perturbation intensity and confidence measures [guo:2024:exploring, liu:2024:effective]. These efforts have contributed to the development of more diversified attack methodologies.
@@ 3.1 Multi-Objective Adversarial Attack
The notation remains consistent with the single-objective scenario as depicted in Equation (1). Furthermore, this paper adopts the extensively utilized $\ell_{\infty}$-ball as the constraint set for perturbations, denoted by $\mathcal{B}=\{\bm{\delta}:\|\bm{\delta}\|_{\infty}\leq\epsilon\}$.
Examples with higher values across multiple surrogate loss functions are more susceptible to misclassification by the model. In existing literature, this statement is supported by frequent misclassifications of adversarial examples with high values on singular loss functions [madry:2018:towards, croce:2020:reliable, yamamura:2022:diversified].
@@ 3.2 Decomposition-Based Optimization
In this research, we employ the Tchebycheff decomposition to transform the multi-objective problem into a suite of single-objective subproblems. Contrary to the linear scalarization method, the Tchebycheff approach is capable of targeting any location on the Pareto front. This is well-recognized in the discipline of multi-objective optimization [deb:2002:fast, zhang:2007:moead, miettinen2012nonlinear]. By adopting this method, given $K$ weight vectors ${\bm{w}}_{1},\ldots,{\bm{w}}_{K}$, the $k$-th decomposed subproblem is defined as follows:
@@
The Tchebycheff method is instrumental in identifying both convex and non-convex parts of the Pareto front with vertical contour lines [wang:2015:pareto], as demonstrated in Figure 1(a). Nonetheless, it presents three challenges in optimization:
  • Complexity: Accurate approximation of the Pareto front necessitates multiple points, exceeding the number of objectives (>m>m).
@@ 3.3 Smooth Set-based Optimization
Smooth Set-based Optimization To smooth the above optimization problem, we need to take advantage of smooth max and smooth min operators [beck:2012:smoothing, lin:2024:smooth, lin:2024:few].
@@
where $\mu$ is a smoothing parameter. A proof of the above approximation can be found in [beck:2012:smoothing].
@@
The above formulation eases the optimization process and avoids oscillations in the optimization process, which is analyzed in the multi-objective literature [lin:2024:smooth]. We illustrate the smoothed set-based optimization problem along with a possible optimization trajectory in Figure 1(c).
@@ 4.1 MOS Attack: Implementation by APGD
By formulating a smooth set-based optimization problem, we can now apply gradient-based optimization algorithms for its efficient resolution. Within the domain of adversarial attacks, our framework incorporates the well-known APGD algorithm [croce:2020:reliable]. In this section, we provide a detailed explanation of our attack, as outlined in Algorithm 1.
@@
0 Cross Entropy Loss [szegedy:2014:intriguing, madry:2018:towards, goodfellow:2015:explaining, kurakin:2017:adversarial, carlini:2017:towards] ${\bm{h}}_{y}({\bm{x}})+\log(\sum_{i=1}^{K}e^{{\bm{h}}_{i}({\bm{x}})})$
1 Marginal Loss [croce:2020:minimally, carlini:2017:towards, gowal:2019:alternative, croce:2020:reliable, sriramanan:2020:guided] ${\bm{h}}_{y}({\bm{x}})+\max_{j\neq y}{\bm{h}}_{j}({\bm{x}})$
2 Difference of Logits Ratio [croce:2020:minimally] $(-{\bm{h}}_{y}({\bm{x}})+\max_{j\neq y}{\bm{h}}_{j}({\bm{x}}))/({\bm{h}}_{\pi_{1}}({\bm{x}})-{\bm{h}}_{\pi_{3}}({\bm{x}}))$
3 Boosted Cross-Entropy Loss [wang:2020:improving] $\log{\bm{p}}_{y}({\bm{x}})-\log(1-\max_{j\neq y}{\bm{p}}_{j}({\bm{x}}))$
4 Searched Loss 1 [xia:2024:tightening] $\sum_{i}\exp(10{\bm{p}}/\max_{j}{\bm{p}}_{j})$
5 Searched Loss 2 [xia:2024:tightening] $\exp(-\max(\text{softmax}({\bm{h}}+2\text{softmax}(5{\bm{h}}))))$
6 Searched Loss 3 [xia:2024:tightening] $\text{softmax}(-\text{softmax}(2\exp({\bm{h}}){\bm{h}}))(\text{softmax}(2{\bm{h}})+2{\bm{y}}_{\text{one-hot}})$
7 Searched Loss 4 [xia:2024:tightening] $(\text{softmax}(\text{softmax}(2{\bm{h}})+{\bm{h}}-{\bm{y}}_{\text{one-hot}})-{\bm{y}}_{\text{one-hot}})^{2}$
@@ 4 Searched Loss 1 [xia:2024:tightening] $\sum_{i}\exp(10{\bm{p}}/\max_{j}{\bm{p}}_{j})$
[…]
@@ 4.2 Automated Synergistic Pattern Mining
Few solutions automatically maximize different loss functions in groups in smooth set-based optimization [lin:2024:few]. To mine these loss synergistic patterns, we propose an automated mining method. This method includes two steps: 1) determining the dominant examples that contribute to the loss maximization, and 2) determining the synergistic pattern of these dominant examples.
@@
CIFAR-10 ($\epsilon=8/255$)
APGD
APGD
ACG
All
(1)*8
MOS-8
MOS-8
Diff.(5)
MOS||CE
@@
0 Rade et al. (2022) [rade:2022:reducing] (ddpm) PreActResNet-18 39.17 39.28 42.45 42.78 (6) 42.59 42.77 +3.49
1 Rade et al. (2022) [rade:2022:reducing] (extra) PreActResNet-18 38.55 38.72 41.63 42.21 (6) 42.03 42.23 +3.51
2 Sehwag et al. (2022) [sehwag:2022:robust] ResNet-18 41.57 41.76 43.10 44.16 (6) 43.79 44.18 +2.42
3 Chen et al. (2020) [chen:2020:adversarial] ResNet-50 45.80 45.95 48.00 48.04 (4) 48.09 48.14 +3.49
4 Gowal et al. (2020) [gowal:2020:uncovering] WideResNet-28-10 34.31 34.46 36.39 36.96 (6) 36.77 36.95 +2.19
5 Wang et al. (2023) [wang:2023:better] WideResNet-28-10 29.72 29.91 31.92 32.44 (6) 32.25 32.49 +2.58
6 Rebuffi et al. (2021) [rebuffi:2021:fixing] WideResNet-28-10 35.97 36.15 38.43 39.05 (6) 38.91 39.14 +2.99
7 Sehwag et al. (2022) [sehwag:2022:robust] WideResNet-34-10 36.85 36.96 38.34 39.36 (5) 38.97 39.38 +2.43
8 Rade et al. (2022) [rade:2022:reducing] WideResNet-34-10 34.29 34.45 36.45 36.97 (6) 36.69 36.94 +2.49
9 Gowal et al. (2021) [gowal:2021:improving] WideResNet-70-16 31.43 31.62 32.54 33.50 (5) 33.33 33.51 +1.89
10 Gowal et al. (2020) [gowal:2020:uncovering] WideResNet-70-16 31.89 32.07 33.34 33.94 (5) 33.72 33.92 +1.85
11 Rebuffi et al. (2021) [rebuffi:2021:fixing] WideResNet-70-16 30.45 30.72 32.41 33.06 (6) 32.79 33.10 +2.38
Average Rank 5.92 4.92 4.00 1.67 3.00 1.33
@@
12 Salman et al. (2020) [salman:2020:do] ResNet-18 70.60 70.74 72.94 74.38 (5) 74.24 74.52 +3.87
13 Salman et al. (2020) [salman:2020:do] ResNet-50 61.38 61.58 62.74 64.92 (7) 64.5 64.94 +3.36
14 Wong et al. (2020) [wong:2020:fast] ResNet-50 70.28 70.46 71.68 73.20 (5) 72.96 73.10 +2.64
15 Engstrom et al. (2019) [engstrom:2019:robustness] ResNet-50 67.62 67.82 67.72 70.12 (5) 69.86 69.92 +2.10
16 Salman et al. (2020) [salman:2020:do] WideResNet-50-2 59.02 59.12 58.92 61.26 (5) 60.76 61.14 +2.02